Solve for the following probabilities (ranges of X values):
a.       P(X ≤ 7) when m = 15
b.      P(9 ≤ X ≤ 18) when m = 15
c.       P(X ≥ 15) when m = 15
d.      P(12 ≤ X < 20) when m = 15

(1) Solve the probabilities on the PROB worksheet. (To four decimal places) The P(X ≤7) when the poisson mean = 15 is ________.Assume a poisson distribution

(2) Solve the probabilities on the PROB worksheet. (To four decimal places) The P(9 ≤ X ≤ 18) when the poisson mean = 15 is ________.Assume a poisson distribution

(3) Solve the probabilities on the PROB worksheet. (To four decimal places) The P(X ≥ 15) when the poisson mean = 15 is________.Assume a poisson distribution

(4) Solve the probabilities on the PROB worksheet. (To four decimal places) The P(12 ≤ X < 20) when the poisson mean = 15 is ________.Assume a poisson distribution

(5) Which of the following Excel formulas can calculate P(10 ≤ X ≤ 20) when the poisson mean = 15 using the CUMULATIVE Distribution function (F(x))?

=poisson(20,15,1)-poisson(9,15,1)

=poisson(21,15,0)-poisson(10,15,0)

=poisson(21,15,1)-poisson(10,15,1)

=poisson(20,15,0)-poisson(9,15,0)

(6) Which of the following Excel formulas can calculate P(X ≥ 13) when the poisson mean = 15 using the PROBABILITY function (p(x))?

=sum(p(100):p(13)), where p(x) = poisson(x,15,0)

=1 - sum(p(13):p(100)), where p(x) = poisson(x,15,0)

=1 - sum(p(0):p(13)), where p(x) = poisson(x,15,0)

=1 - sum(p(0):p(12)), where p(x) = poisson(x,15,0)
Answer #1

The poisson probability is given as:

P(X=x)=\frac{e^{-m}m^x}{x!}, x=0,1,2,.....\infty

1)

P(X\leq 7)=P(X=0)+P(X=1)+....+P(X=7)

P(X=0) 3.05902E-07
P(X=1) 4.58853E-06
P(X=2) 3.4414E-05
P(X=3) 0.00017207
P(X=4) 0.000645263
P(X=5) 0.001935788
P(X=6) 0.00483947
P(X=7) 0.010370294
Total= 0.018002193

P(X < 7) 0.0180

Excel formula used :
Poisson(x,15,0) where x =0,1,.....7
and then summing all the probabilities
Alternate formula:
Poisson(7,15,1) will give the direct result

2)

P(X 18)

P(X=9) 0.032407
P(X=10) 0.048611
P(X=11) 0.066287
P(X=12) 0.082859
P(X=13) 0.095607
P(X=14) 0.102436
P(X=15) 0.102436
P(X=16) 0.096034
P(X=17) 0.084736
P(X=18) 0.070613
Total 0.782025

P(9 < X < 18)-0.7820

Excel formula used :
Poisson(x,15,0) where x =9,10,.....18
and then summing all the probabilities
Alternate formula:
Poisson(18,15,1)-Poisson(8,15,1) will give the direct result.

3)

P(X 15)1- P(X <15) 1- P(X < 14) |-|P(X = 0) + P(X = 1) + + P(X = 14)1

P(X=0) 3.06E-07
P(X=1) 4.59E-06
P(X=2) 3.44E-05
P(X=3) 0.000172
P(X=4) 0.000645
P(X=5) 0.001936
P(X=6) 0.004839
P(X=7) 0.01037
P(X=8) 0.019444
P(X=9) 0.032407
P(X=10) 0.048611
P(X=11) 0.066287
P(X=12) 0.082859
P(X=13) 0.095607
P(X=14) 0.102436
Total= 0.465654

PlX > 15) = 1-P(X < 15) = 1-P(X < 14) 1-0.4657 = 0.5343

Excel formula used to calculate probability upto 14 :
Poisson(x,15,0) where x =0,1,.....14
and then summing all the probabilities and subtracting from 1

Alternate formula:
1-Poisson(14,15,1) will give the direct result

4)

P(12 < X < 20) = P(X 12) + PlX-13) + P(X-20)

12 P(X=12) 0.082859
13 P(X=13) 0.095607
14 P(X=14) 0.102436
15 P(X=15) 0.102436
16 P(X=16) 0.096034
17 P(X=17) 0.084736
18 P(X=18) 0.070613
19 P(X=19) 0.055747
20 P(X=20) 0.04181
Total 0.732277

P(12 < X < 20)-0.7323

Excel formula used :
Poisson(x,15,0) where x =12,13,.....20
and then summing all the probabilities
Alternate formula:
Poisson(20,15,1)-Poisson(11,15,1) will give the direct result.

(5) Which of the following Excel formulas can calculate P(10 ≤ X ≤ 20) when the poisson mean = 15 using the CUMULATIVE Distribution function (F(x))?
=poisson(20,15,1)-poisson(9,15,1)

(6) Which of the following Excel formulas can calculate P(X ≥ 13) when the poisson mean = 15 using the PROBABILITY function (p(x))?

=1 - sum(p(0):p(12)), where p(x) = poisson(x,15,0)